OpenStudy (anonymous):

Hey. Guys please help (5a-2b)(a^2-ab-b^2)

OpenStudy (anonymous):

Multiply verticaly

OpenStudy (marigirl):

5a^3-7a^2b-3ab^2+2b^3

OpenStudy (marigirl):

is that what u meant?

OpenStudy (anonymous):

Yes I think so

OpenStudy (anonymous):

Do u also happen to know about sequence of numbers

OpenStudy (marigirl):

like finding rules ? sequence and series?

OpenStudy (marigirl):

i could try

OpenStudy (anonymous):

Ok sweet find the next number in this sequence. 1,8,16,25

OpenStudy (marigirl):

so u just need to nxt number? I believe it will be 35

OpenStudy (anonymous):

How did u get that. I got something else

OpenStudy (anonymous):

8-1 = 7 ,16-8 =8 ,25-16 = 9,, so x -25 = n+1 you should be able to see the difference in pattern by 7 then 8 then 9 then so it's n +1

OpenStudy (marigirl):

yes, with a difference of 10 then? 1 8 16 25 35 7 8 9 10 Difference